why are E1 and E2 subtracted from each other for Enet?
Answer #1

If E1 and E2 are opposite in direction at a point, to find magnitude of net electric field we subtract E1 from E2 (or) E2 from E1.

If E1 is smaller than E2, |Enet| = E2 - E1

If E1 is greater than E2, |Enet| = E1 - E2

If E1 and E2 both have same direction, |Enet| = E1 + E2